THE ROTATIONAL SPECTRUM OF AMINOETHANOL (NH_{2}CH_{2}CH_{2}OH)

\underline{SUSANNA L. WIDICUS}, Division of Chemistry, California Institute of Technology, Pasadena, CA 91125; BRIAN J. DROUIN, California Institute of Technology, Jet Propulsion Laboratory, Microwave Millimeter and Submillimeter Spectroscopy Lab, Pasadena, CA 91109; KATHRYN A. DYL, and GE-OFFREY A. BLAKE, Division of Geology and Planetary Science, California Institute of Technology, Pasadena, CA 91125.

The rotational spectrum of aminoethanol (NH_{2}CH_{2}CH_{2}OH) has been characterized in the 75 – 120 GHz region using the 1-meter flow cell apparatus at the Jet Propulsion Laboratory. Previous studies of the rotational spectrum of aminoethanol in the 8 – 40 GHz region probed the ground state to the J = 8 level and included determination of the dipole moment and rotational constants.^{a} Fourth-order centrifugal distortion constants were subsequently determined from these data.^{b} We present here extended measurements of the ground state transitions up to J = 51, as well as ν_{27}=1 transitions up to J = 50, ν_{26}=1 transitions up to J = 35, and ν_{25}=1 transitions up to J = 29. Rotational transitions from the ν_{27}=2, ν_{26}=2, and ν_{25}=2 states as well as three combination bands have also been assigned. Spectroscopic parameters, including standard rotational constants and fourth- and sixth-order centrifugal distortion constants, were determined using the SPFit/SPCat program suite.

aPenn R. E. and Curl R. F. (1971) J. Chem. Phys. 53, 651 - 658.

bKaushik V. K. and Woods R. C. (1982) Z. Phys. Chem. Neue Fol. 132, 117 – 120.